How much energy (heat) is required to convert 52.0 gram of ice at 10.0C to steam at 100C?
1
Expert's answer
2021-10-12T07:48:23-0400

You need to break it down. Energy when not in a phase change is Q=mc(deltaT) whereas energy at a phase change is Q=deltaH(m)


So from -10C to 0C, no phase change occurs and it is ice:

Q=52g*2.09 J/g C*(10C)= 1086.8J=1.087kJ


At 0C, melting occurs, so you use deltaH fusion:

Make sure your units work out properly:

Q=6.02kJ/mol(52g)(1mol/18g) = 17.39 kJ


From 0C to 100C, no phase change occurs when it's liquid:

Q=52g*4.18 J/g C*(100C)= 21736J=21.736 kJ


At 100C, phase change occurs as water evaporates:

Q=40.7kJ/mol*52g*(1mol/18g) = 117.6 kJ


Add these values of heat up: 1.087 + 17.39+21.736+117.6=157.8 kJ